# Privacy for Quantum Annealing. Attack on Spin Reversal Transformations in the case of cryptanalysis

### v1

@@ 3.3 Executing the attack, the most straightforward phase
Using a parameterized problem, an attack is performed. A vector of biases is sufficient to perform the attack. The resulting encrypted vector of biases is juxtaposed with a parameterized one to form a system of equations. As can be seen from the construction of the parameterized matrix, each of the equations of the system can be presented as Equation (21). In general, the coefficients of a vector can be the sum of a rational number and a linear combination of any keystream bits, as $\mathcal{A}_{i}$, we denote the set of indices of the relevant bits.
$h_{i}^{*}=\sum_{j\in\mathcal{A}_{i}}+a_{i},$ (21)
where $a_{i}\in\mathbb{Q}$ The next step is to analyze each of the resulting equations in the correct order. First, the equations must be transformed to form presented as Equation (22)
$z_{j}+b_{i}=h_{i}^{*},$ (22)
where, $b_{i}\in\mathbb{Q}$. If the transformation of the equations is not possible, reduce the system to upper triangular form. For example, using the Gaussian reduction. Then, an equation analysis is performed from the last equation, considering previous solutions. The variable $z_{j}$ is a binary variable, $z_{j}\in\{0,1\}$, therefore the relation (23) should be satisfied.
@@
If the above relation is satisfied, then:
@@
  • designated bit of concealment key is: $x_{i}=0$;
  • designated bit of keystream is: $z_{j}=h_{i}^{*}-b_{i}$.
@@
  • designated bit of concealment key is: $x_{i}=1$;
  • designated bit of keystream is: $z_{j}=-h_{i}^{*}-b_{i}$.
In most cases, solving the equation above gives the bit of the keystream $z_{j}$ with a probability of $1$. If one finds that $z_{j}\in\{0,1\}$ regardless of whether a coefficient is concealed, one can check if other equations where $z_{j}$ appears are correct. The above procedure allows the recovery of the used keystream. The remaining bits of the concealment key can be determined by comparing the remaining coefficients. If the coefficients are opposite, the given coefficient has been concealed, and the corresponding concealment key bit is 1.
For every bit of the keystream, retrieving the correct value of the bit requires only solving an affine equation, and the whole attack is swift. In addition, the attack allows the recovery of the concealment key. With the use of a key:
@@
  • the solution to the problem sent may be unveiled, including knowledge of the secret key, which the client wanted to recover.
@@ 4 A practical example of the proposed attack
To illustrate the correctness of the attack, an attack performed in practice is presented. A scaled-down $E_{0}$ cipher was used to show the step-by-step operation of the attack. In addition, the execution of the attack on the full version was discussed.
@@ 4.1 An illustrative attack on a scaled-down instance
It should be emphasized that to simplify the generated problem as much as possible the registers of the scaled cipher are of equal length. For the selected cipher instance, the following keystream was used:
@@
For the given sequence, 96 equations were generated. Based on the equations, a QUBO problem with 240 variables was determined. The problem was transformed to an Ising model and encrypted using the idea presented in [8]. The relevant part of the used key is presented as equation (25).
@@
An oracle was built based on [21]. Using it, a parameterized problem was generated according to the idea in the section 3.2. From the parameterized vector of biases, the equations presented as Equation (26) are selected to recover the keystream. Out of 240 equations, 12 equations were selected for further analysis.
@@
Finally, based on the above analysis, the keystream was determined. The entire concealment key was recovered using the remaining constant coefficients of the matrix. The key is shown in the Figure 5.
@@
Missing bits, denoted as $x$ in the Figure 5, of the key do not affect the ciphertext, it is not possible to change the sign of the coefficient equal to 0. A comparison of the recovered keystream and the key confirms the correctness of the attack.
@@ 4.2 Attack on the full version of the $E_{0}$ cipher
An attack on an encrypted problem corresponding to the cryptanalysis of the complete cipher $E_{0}$ can be performed similarly. The problem has 3071 variables, and the wanted keystream has 129 bits. As in the rescaled example, the longest step is the parameterization of the matrix. The time required to generate the parameterized matrix may be estimated for about 13.5 hours. It should be noted that the script used is not optimized in any way, and the experiment only illustrates the disparity between the time of the actual phase of the attack and the time of the pre-computations. The attack phase requires solving only 129 affine equations of one variable.
As can be seen from the above time disparity, once the longer pre-calculations are performed once, the time required to execute the attack is practically imperceptible.
@@ 5 Summary and future work
This paper shows a practical attack on a proposed scheme to ensure the privacy of problems sent to quantum computing clouds. In addition, a practical attack on a small version of the cipher was demonstrated to show the correctness of the attack. It should be noted that the work does not address the correctness of the scheme under attack. The proposed attack was prepared with assumptions in the original work and additional ones resulting from the functioning of available services.
Further work should look for a secure way to secure cloud computing based on the Ising model. As an alternative to using an untrusted provider, methods can be developed that allow local, using private infrastructure, to solve the problems discussed.

### v3

@@ 4.4 Executing the attack, the most straightforward phase
An attack is performed using a parameterized problem. A vector of biases is sufficient to perform the attack.
First, a system of linear equations is constructed. The encrypted vector of biases $h^{*}$ is juxtaposed with a parameterized one $h^{P}$ to form a system of equations:
$h^{*}=h^{P}=\begin{bmatrix}h^{*}_{0}\\ h^{*}_{1}\\ \vdots\\ h^{*}_{N}\end{bmatrix}=\begin{bmatrix}h^{P}_{0}\\ h^{P}_{1}\\ \vdots\\ h^{P}_{N}\end{bmatrix}.$
As the design of the parameterized matrix shows, each element of $h^{P}$ can be presented as a sum of rational number $\texttt{const}_{i}$ and a linear combination of selected bits of the keystream. We denote the set of indices of the relevant bits for element $i$ as $\mathcal{A}_{i}$. The equations formed in the proposed way can be represented by Equation (17):
$h_{i}^{*}=\sum_{j\in\mathcal{A}_{i}}z_{j}+\texttt{const}_{i},$ (17)
where $\texttt{const}_{i}\in\mathbb{Q}$.
The next step is to analyze each resulting equation in the correct order. According to how the problem is constructed, there will be more equations than unknowns in the created matrix. From among all the equations, we choose some set of sufficient size and reduce the equations to the form presented in Equation (18):
$z_{k}+b_{i}=h_{i}^{*},$ (18)
where $b_{i}\in\mathbb{Q}$ is sum of $\texttt{const}_{i}$ and known keystream bits. As $k$ we denotes the index of the keystream variable occurring in this equation.
If the transformation of all the equations is impossible, we rearrange the system to an upper triangular form. Then, an equation analysis is performed, starting with the last equation. In the next steps, successive equations will be reduced, considering the previous solutions and performing an analysis.
As can be seen from the definition of the optimization problem under consideration, the variable $z_{k}$ is a binary variable, $z_{k}\in\{0,1\}$. Therefore, the relation (19) should be satisfied:
@@
Due to the use of encryption, two situations can occur: the condition will be met or not. If the above relation is satisfied, then:
@@
  • the designated bit of the concealment key is $x_{i}=0$;
  • the designated bit of the keystream is $z_{k}=h_{i}^{*}-b_{i}$.
@@
  • the designated bit of the concealment key is $x_{i}=1$;
  • the designated bit of the keystream is $z_{k}=-h_{i}^{*}-b_{i}$.
In most cases, solving the equation above gives the bit of the keystream $z_{k}$ with a probability of $1$. If one finds that $z_{k}\in\{0,1\}$ regardless of whether a coefficient is concealed, one can check if other equations where $z_{k}$ appears are correct.
The above procedure allows the recovery of the used keystream. The remaining bits of the concealment key can be determined by comparing the remaining coefficients. If the coefficients are opposite, the given coefficient has been concealed, and the corresponding concealment key bit is 1. For every bit of the keystream, retrieving the correct value of the bit requires only solving an affine equation, and the entire attack is swift. In addition, the attack allows the recovery of the concealment key. With the use of this key:
@@
  • the solution to the problem sent may be unveiled, including knowledge of the secret key that the client wanted to recover.
@@ 5 A practical example of the proposed attack
To illustrate the correctness of the attack, an attack performed in practice is presented. A scaled-down $E_{0}$ cipher was used to show the step-by-step operation of the attack. In addition, the execution of the attack on the full version is discussed.
5.1 Example parameterization
In this section, we present a pen-and-paper example of parameterization. We are given an oracle $\phi$ that returns bias vectors for given streams generated by a hypothetical stream cipher. Here, we focus on supporting the idea of parameterization with an example, so we omit the details of the hypothetical cipher. Assume that we are looking for a parameterized matrix corresponding to an algebraic attack using a 3-bit keystream $(z_{0},z_{1},z_{2})$. We now follow the description presented earlier:
  1. 1.
    We determine the bias vector for a stream with a Hamming weight of 0:
    $h_{\infty}=\phi(0,0,0)=\begin{bmatrix}2\\ 4\\ 1\end{bmatrix}.$
  2. 2.
    We determine the bias vector for all streams with a Hamming weight of 1:
    $\begin{array}[]{ccc}h_{0}=\phi(1,0,0)=\begin{bmatrix}3\\ 5\\ 2\end{bmatrix},&h_{1}=\phi(0,1,0)=\begin{bmatrix}2\\ 5\\ 1\end{bmatrix},&h_{2}=\phi(0,0,1)=\begin{bmatrix}2\\ 4\\ 7\end{bmatrix}.\end{array}$
  3. 3.
    We determine the differences $h_{i}-h_{\infty}$, where $i\in\{0,1,2\}$:
    $\begin{array}[]{ccc}h_{0}-h_{\infty}=\begin{bmatrix}1\\ 1\\ 1\end{bmatrix},&h_{1}-h_{\infty}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix},&h_{2}-h_{\infty}=\begin{bmatrix}0\\ 0\\ 6\end{bmatrix}.\end{array}$
  4. 4.
    We determine the parameterized vector as:
    $h^{P}=h_{\infty}+(h_{0}-h_{\infty})\cdot\texttt{z}_{\texttt{0}}+(h_{1}-h_{\infty})\cdot\texttt{z}_{\texttt{1}}+(h_{2}-h_{\infty})\cdot\texttt{z}_{\texttt{2}},$
    so:
    $h^{P}=\begin{bmatrix}2\\ 4\\ 1\end{bmatrix}+\begin{bmatrix}\texttt{z}_{\texttt{0}}\\ \texttt{z}_{\texttt{0}}\\ \texttt{z}_{\texttt{0}}\end{bmatrix}+\begin{bmatrix}0\\ \texttt{z}_{\texttt{1}}\\ 0\end{bmatrix}+\begin{bmatrix}0\\ 0\\ 6\texttt{z}_{\texttt{2}}\end{bmatrix}=\begin{bmatrix}2+\texttt{z}_{\texttt{0}}\\ 4+\texttt{z}_{\texttt{0}}+\texttt{z}_{\texttt{1}}\\ 1+\texttt{z}_{\texttt{0}}+6\texttt{z}_{\texttt{2}}\end{bmatrix}.$
The searched parameterized matrix of the assumed algebraic attack was thus determined using four oracle queries.
@@ 5.2 An illustrative attack on a scaled-down instance
To simplify the generated problem as much as possible, the registers of the scaled cipher should be of equal length. The set of primitive polynomials of degree 3 is limited. For this reason, the selected polynomials are equal. The impact of such a solution on the cipher’s security is not the subject of this paper. Such a configuration allows a pictorial representation of the proposed attack.
For the selected cipher instance, the following keystream was used:
@@
For the given sequence, 96 equations were generated. Based on these equations, a QUBO problem with 240 variables was determined. The problem was transformed into an Ising model and encrypted using the idea presented in [9], as described in Section 3. Due to the size of the matrix, we cannot include the entire key here. Below, as Equation (21), we present selected key bits corresponding to the coefficients of the vector at the positions analyzed:
@@
Values under consecutive bits indicate the position number of the specified key bit.
An oracle was built based on [26]. Using it, a parameterized problem was generated according to the idea in Section 4.3. To obtain the entire parameterized bias vector, 13 queries were performed. The equations presented as Equation (22) are selected to recover the keystream from this vector. As described in Section 4.4, such a set of equations was chosen so that it would be possible to analyze them one by one and determine the entire keystream. Out of 240 equations in a parameterized matrix, 12 were selected for further analysis:
@@
Finally, based on the above analysis, the keystream was determined.
The relevant part of the concealment key was determined using the remaining elements of the vector. The recovered key, the result of the developed script implementing the attack, is shown in Figure 2. Missing bits, denoted as $x$ in Figure 2, of the key do not affect the ciphertext. It is impossible to change the sign of a coefficient equal to 0.
@@
We can verify the correctness in two ways:
  • performing a comparison of the sequence shown in Equation (21) with the last column of Table 1;
  • using the recovered keystream, the full bias vector is determined. Using the recovered key, the concealed vector is decrypted. Then, a comparison of both vectors is performed.
Both of these methods confirm the correctness of the attack performed.
@@ 5.3 Attack on the full version of the $E_{0}$ cipher
An attack on an encrypted problem corresponding to the cryptanalysis of the full $E_{0}$ cipher can be performed similarly. Using a script in the SageMath environment:
  1. 1.
    A problem corresponding to an algebraic attack on the $E_{0}$ cipher is generated for the selected keystream. The problem has 2728 variables, and the used keystream has 128 bits.
  2. 2.
    The resulting optimization problem is encrypted using Spin Reversal Transformation with a randomly generated key.
  3. 3.
    A parameterized matrix is determined. It requires 129 oracle calls. The time needed to generate the parameterized matrix can be estimated at about 13.5 hours.
  4. 4.
    Based on the determined matrix, a system of 128 linear equations is arranged.
  5. 5.
    In the last step, the designated equations are analyzed one by one, and the key used and the stream for which the problem was generated is determined.
  6. 6.
    Finally, the results obtained are verified. The recovered keystream is compared with the original one. The matrix is decrypted and compared with the problem generated for the recovered keystream.
As in the rescaled example, the longest step is the parameterization of the matrix. However, the duration of the attack still allows it to be performed in practice.
It should be noted that the script used is not optimized in any way, and the experiment only illustrates the disparity between the time of the actual phase of the attack and the time of the pre-computations.
@@ 6 Summary and future work
This paper shows a practical attack on a proposed scheme to ensure the privacy of problems sent to quantum computing clouds. In addition, a practical attack on a miniature version of the cipher was demonstrated to show the correctness of the attack. It should be noted that the work does not address the correctness of the scheme under attack. The proposed attack was prepared with assumptions in the original work and additional ones resulting from the functioning of available services.
Further work should look for a secure way to protect cloud computing based on the Ising model. As an alternative to using an untrusted provider, methods can be developed that allow local problem solving using private infrastructure.